OpenStudy (anonymous):

What do you get when you solve 3 radical 13 plus 7 radical 13

OpenStudy (anonymous):

here is the equation: 

OpenStudy (saifoo.khan):

\[3 \sqrt{13} + 7 \sqrt{13}\] \[(3+ 7) \sqrt{13}\] \[10 \sqrt { 13}\]

OpenStudy (anonymous):

thanks! :)

OpenStudy (saifoo.khan):

welks! :D

OpenStudy (anonymous):

You keep the radical the same bcuase its common to both terms and a the coeficients so add7 and 3 to gt10 radical 13

OpenStudy (anonymous):

thanks appreciate it! :D